2. If the price of good B increases by 2.6% and the price elasticity of demand is 0.5, find the percentage change in quantity demanded and the percentage change in revenue. If you want to increase revenue should you increase or decrease the price in this case?
Answer #1

Ans) percentage change in quantity demanded = elasticity * percentage change in price

= -0.5*2.6 = -1.3 ( decrease)

Percentage change in revenue = percentage change in price - percentage change in quantity demanded

= 2.6 - 1.3 = 1.3 ( increase)

If you want to increase revenue you should increase the price in this case because the demand is inelastic